OpenStudy (anonymous):

solve (2a-1)(4a5)/(2a-3)(4a-5) - (2a-4)(2a-1)/(2a-3)(4a-5) - (3a-2)(2a-3)/(2a-3)(4a-5)

OpenStudy (anonymous):

saying "solve" implies there is an equation, this is an expression

OpenStudy (anonymous):

no.... its wrote as fractions.. I was trying to draw it out to show how it looks but it's too much to draw :o

OpenStudy (anonymous):

\[\frac{(4 a-5) \left(a \left(8 a^5-4 a^4-10 a+23\right)-10\right)}{2 a-3} \]